Grade 11Algebra

A cylinder contains acetylene gas at 27 C and 4.04M Pa.the pressure ​in the cylinder after half the mass if gas is used up and temperature has fallen to 12 C will be what??.....plz solve this with complete steps and formula ...and plz explain why cant we use the formula. P1/T1 = P2/T2 for solving the problem???

Avlesh Patel
8 Years agoGrade 11

1 Answer

Eshan
ApprovedApproved Tutor Answer8 Years ago
Dear student,

Ideal gas law is always valid. HencePV=nRTholds true. If in a particular situation it is given that both the volume and number of moles of gas remains constant(that is,V,nare constants), you may use the relation\dfrac{P}{T}= constant.

However in this case it is given that half the mass of gas is used up, and hence half of the number of moles of gas remain, but the volume remains constant.

Hence\dfrac{P}{nT}=constant\implies \dfrac{P_1}{n_1T_1}=\dfrac{P_2}{n_2T_2}

\implies \dfrac{4.04MPa}{n\times 300K}=\dfrac{P_2}{\dfrac{n}{2}\times 285}=1.92MPa